# Complements of discriminants of real parabolic function singularities. II

### v2

@@ 2.3. Enumeration of components for $X_{9}^{+}$ singularities
Pictures $(f)$ and $(g)$ of Fig. 1 are topologically equivalent. However, Bezout’s theorem arguments show that the polynomials that realize them cannot be connected within the space of non-discriminant polynomials of class $\Phi_{+}$. Figure 1 coincides with the well-known list of affine rigid isotopy classes of bounded degree four smooth algebraic plane curves. In our case, however, the isotopies are allowed only within a much smaller space $\Phi_{+}$. In other cases, $\Phi_{1}$ and $\Phi_{2}$, the analogous restriction leads to an increase of classification lists.
@@
Using the Lyashko–Looijenga covering (in the form described in [10]), we can assume that the functions $f$ and $\tilde{f}$ have the same set of critical values and the same set of paths in ${\mathbb{C}}^{1}$ specifying the common associated virtual function. Let $I:[0,1]\to\Phi_{+}$ be a generic path such that $I(0)=f$, $I(1)$ is the polynomial ${\bf T}(x^{4}+x^{3}+y^{4})$, and the path $I$ has in its interior points only finitely many transversal intersections with the variety of non-generic polynomials in $\Phi_{+}$ (that is, with the union of the discriminant and the set of not strictly Morse functions). Since the virtual function of $\tilde{f}$ is the same as of $f$, there exists another path $\tilde{I}:[0,1]\to\Phi_{+}$ such that $\tilde{I}(0)=\tilde{f}$, the polynomials $I(t)$ and $\tilde{I}(t)$ have the same set of critical values for any $t\in[0,1]$, and the polynomials of the path $\tilde{I}$ undergo the surgeries of the same type as $I$ at the same instants $t$. (For this existence statement, the parabolicity of $X+9$ singularities is essential.) In particular, the point $\tilde{I}(1)$ can only be one of the four points (2). First, suppose that it is the same point $I(1)$. A neighborhood of this point in the space $\Phi_{+}$ has only one nontrivial smooth automorphism that preserves the associated virtual functions and commutes with the Lyashko–Looijenga map: this is the map induced by the reflection of ${\mathbb{R}}^{2}$ in the line $y=0$. Therefore, there are only two possibilities: either the final segments of the paths $I$ and $\tilde{I}$ (for $t\in(1-\varepsilon,1]$) are the same, or they are mapped into each other by this automorphism. In both cases, this condition can be continued to the entire paths $I$ and $\tilde{I}$ and proves that either $\tilde{f}=f$ or these two functions are mapped to each other by this automorphism, which belongs to the group $G_{0}$. If the point $\tilde{I}(1)$ is different from $I(1)$, we first apply the element of the group $G_{0}$, moving these points to each other, to the entire path $\tilde{I}$ (in particular, to its starting point $\tilde{f}$). This leads us to the situation considered previously. $\Box$
Proof of Theorem 5. By Proposition 8, any virtual component mentioned in Proposition 7 is associated with at most eight real connected components. Let us prove that in fact each virtual component represented by a polynomial shown in Fig. 1 is associated with a single real connected component.
@@
with arbitrary $A\in(-1,1)$. This polynomial is invariant under the transformations $(x,y)\mapsto(x,-y)$ and $(x,y)\mapsto(-x,y)$. Thus, its $G_{0}$-orbit consists only of itself and the polynomial
@@
Proof. If we replace $A$ by 1 in (5), then all these polynomials will be obtained from each other by the rotations of ${\mathbb{R}}^{2}$ by angles $\varphi\in[0,\pi/2]$. Each of these polynomials has only three real Morse critical points: one at the origin and two on the unit circle. If $A=1-\varepsilon$ with sufficiently small $\varepsilon$, then the real zero-level set deforms only slightly in the domain $\{x^{2}+y^{2}\leq 2\}.$ In particular, it maintains its topological structure in this domain for any $\varphi$. It also remains empty outside this domain. $\Box$
@@ 2.4. The 1-cohomology of discriminant complements of $X_{9}^{+}$ singularities
Recall the notation $B({\mathbb{R}}^{2},k)$ for the unordered configuration space, that is, the space of $k$-point subsets of the plane. For any $k\geq 2$, $H^{1}(B({\mathbb{R}}^{2},k))\simeq{\mathbb{Z}}$. The basic element of this group is induced from $H^{1}({\mathbb{C}}^{1}\setminus\{0\})$ by the map $B({\mathbb{R}}^{2},k)\equiv B({\mathbb{C}}^{1},k)\to{\mathbb{C}}^{1}\{0\}$ sending any collection of $k$ distinct point $z_{1},\dots,z_{k}$ to the product of all k(k1)k(k-1) complex numbers $z_{i}-z_{j}$, $i\neq j$. Let $U$ be a connected component of the discriminant complement in $\Phi_{+}$ that consists of polynomials whose zero-level sets are $k\geq 2$ non-nested ovals. Then a map $H^{*}(B({\mathbb{R}}^{2},k))\to H^{*}(U)$ is defined as follows. For any polynomial $f\in U,$ define $\beta(f)$ as the space of all choices of a point inside any of its ovals. These spaces $\beta(f)$ form a locally trivial fiber bundle over the component $U$. The fibers of this fiber bundle are contractible, hence it has a cross-section, which is a continuous map $U\to B({\mathbb{R}}^{2},k)$. Clearly, these maps defined by different cross-sections are homotopic and define the same homomorphism of cohomology rings.
@@ 2.5. The components for the $X_{9}^{1}$ singularities
acts on the space $\Phi_{1}$. It preserves the set of discriminant polynomials and hence permutes the connected components of its complement.
This involution naturally extends to the set of associated virtual components: each virtual component which is the value of the set-valued invariant of the polynomial $f(x,y)$ goes to the value of the polynomial f(x,y)-f(x,-y). These involutions have no fixed elements: indeed, the sum of the values of the invariant Ind (see Definition 3) for any two components (real or virtual) related by the involution is always equal to 1-1.
@@
There are exactly ten virtual components of type $X_{9}^{1}$. They split into five pairs of virtual components related to each other by the involution induced by ((13)). The topological types of zero-level sets of polynomials associated with elements of some five of the virtual components ((one from each pair)) are shown in pictures ((a)), ((b)), ((c)), ((d), and ((f)) of Fig. 2. $\Box$
@@
Proof. Let $f$ be a polynomial that satisfies the conditions of this lemma. By Lemma 5 of [21], the $j$-invariant of its principal homogeneous part is equal to 0, hence $A=\pm\frac{\sqrt{3}}{2}$ in (12). In the affine coordinates $(x,y)$ centered at the $E_{7}$ point, the monomials of degree $\leq 2$ of this polynomial vanish, and the homogeneous part of degree three has the form $\alpha x^{3}$ or $\alpha y^{3}$, $\alpha\neq 0$. The critical value of this function at the $A_{2}$ point is equal to $\pm 3^{\frac{15}{2}}2^{-8}\alpha^{2}$, where $\pm$ is the sign of $A$. This value is equal to $3^{\frac{15}{2}}2^{-8}$ if and only if this sign is ++ and $\alpha$ equals $1$ or 1-1. $\Box$
Proof of Proposition 15. The deduction of this proof from Lemma 16 essentially repeats the proof of Proposition 8 with the following simplification. In this case, the neighborhood of the point ${\bf\tilde{T}}(x^{3}y+\sqrt{3}x^{2}y^{2}+xy^{3}+x^{3})$ in the parameter space of the deformation (1) has no nontrivial involutions that commute with the Lyashko–Looijenga covering (see [21], § 4.5). Therefore, the statement of Proposition 15 follows immediately from the fact that all four polynomials (14) belong to the same orbit of the group $G_{1}$. $\Box$
@@
Picture $(d)$ (respectively, $(e)$) is realized by the polynomial
(15) ${\bf\tilde{T}}\left((x^{2}+y^{2}-t^{2})(x+0.9t)(y+0.9t)+\alpha t^{4}\right),$
$\alpha\in(0,0.003)$, with positive (respectively, negative) $t$. These two pictures are topologically equivalent. However, it easily follows from Bezout’s theorem that these polynomials represent different connected components of the discriminant complement in the space $\Phi_{1},$ see [21].
@@ 2.6. Components for the $X_{9}^{2}$ singularities
There are exactly eight polynomials ${\mathbb{R}}^{2}\to{\mathbb{R}}$ of class $\Phi_{2}$ that have a critical point of class $D_{6}$ with critical value 0 and a critical point of class $A_{3}$ with critical value $\frac{-1}{324}$. These eight polynomials form an orbit of the natural action of the group $G_{0}$.
@@
The rest of the proof of Proposition 20 repeats the proof of Proposition 3 of [21] in § 5.4 of [21]. $\Box$
@@
whose zero-level set is topologically equivalent to the picture (d), but has a singular point of class $E_{6}$ at its center. Then, we perturb this singular point by adding very small terms of lower degrees. These terms do not change the topology of the curve far away from this singular point, but provide the standard dissolutions of the $E_{6}$ singularity, see Figure 24 in [24]. All the resulting polynomials (2630) are invariant under the reflection in the axis $y=0$. Thus, in each of these five cases, the virtual component is associated with no more than four real connected components of the discriminant complement. These four components differ for the topological reasons.
@@ 3. Components for $J_{10}$ singularities
Denote by $\Omega_{1}$ and $\Omega_{3}$ the spaces of all polynomials of the form (31) and (32), respectively. The formal graph and its virtual components for $J_{10}$ singularities are defined in the same way as for the case $X_{9}$, except that the frames of vanishing cycles consists of 10 elements.
@@
If the virtual components of two generic polynomials $f$, $\tilde{f}$ of class $\Omega_{1}$ or $\Omega_{3}$ are the same, then their connected components of the space of generic polynomials either are the same or are mapped to each other by the involution
@@
In particular, any virtual component of class $\Omega_{1}$ or $\Omega_{3}$ is associated with at most two real connected components. $\Box$
In the following pictures, the $x$ coordinate axis is directed upwards, and the horizontal line is the axis $\{x=0\}$.
@@ 3.2. Components for $J_{10}^{1}$ singularities
There are exactly ten virtual components of the formal graph of type $J_{10}^{1}$. These virtual components are represented by polynomials, the topological shapes of whose zero-level sets are shown on Fig. 4. $\Box$
@@
There are exactly thirteen connected components of the discriminant complement in the space of all polynomials of class $\Omega_{1}$. Each of the virtual components shown in pictures $(a)$, $(b)$, (b~)($\tilde${b}), $(c)$, (c~)($\tilde${c}), $(d)$, and (d~)($\tilde${d}) of Fig. 4 is associated with a single real connected component. Each of the virtual components/topological types represented in pictures $(e)$, (e~),($\tilde${e}), and $(f)$ is associated with two different components that are related by the involution ((33)).
@@ 3.3. Components for $J_{10}^{3}$ singularities
(35) (h),(h~),(i),(i~),(j),(j~),(k),(k~),(g),(l)(h),\ ($\tilde${h}),\ (i),\ ($\tilde${i}),\ (j),\ ($\tilde${j}),\ (k),\ ($\tilde${k}),\ (g),\ (l)
@@ 4. The cases $P_{8}$
with $A<1$ for $P_{8}^{1}$ singularities and $A>1$ for $P_{8}^{2}$. The zero sets of these normal forms of class $P_{8}^{1}$ (respectively, $P_{8}^{2}$) define cubic curves with one (respectively, two) connected components in ${\mathbb{R}P}^{2}$. Thus, each polynomial of class $P_{8}^{1}$ defines a conical surface in ${\mathbb{R}}^{3}$ that is homeomorphic to the plane with a single singular point at the origin. The zero set of a polynomial of class $P_{8}^{2}$ consists of such a surface and a surface that is homeomorphic to the standard quadratic cone.
It follows easily from Morse theory that the closure in ${\mathbb{R}P}^{3}$ of the set of zeros of any non-discriminant perturbation $f_{\lambda}$ of a singularity (41) with $\mbox{\rm Ind}(f_{\lambda})$ equal to 0 ((respectively, 1-1, 2-2, 3-3, or 1)) is homeomorphic to the projective plane ((respectively, the projective plane with one, two, three handles, or the disjoint union of the projective plane and the sphere)).
@@
A polynomial $f:({\mathbb{C}}^{3},{\mathbb{R}}^{3})\to({\mathbb{C}},{\mathbb{R}})$ of the form (42) is generic if it has only Morse critical points in ${\mathbb{C}}^{3}$, its critical values are all different and not equal to 0, and the complex zero-level set of the principal homogeneous part of $f$ defines a smooth curve in ${\mathbb{C}}P^{2}$.
Let $\Xi_{1}$ and $\Xi_{2}$ denote the parameter spaces ${\mathbb{R}}^{7}\times(-\infty,1)$ and ${\mathbb{R}}^{7}\times(1,\infty)$ of the deformations (42) of the functions (41) of classes $P_{8}^{1}$ and $P_{8}^{2}$.

### v3

@@ 2.3. Enumeration of components for $X_{9}^{+}$ singularities
Pictures $(f)$ and $(g)$ of Fig. 1 are topologically equivalent. However, Bezout’s theorem arguments show that the polynomials that realize them cannot be connected within the space of non-discriminant polynomials of class $\Phi_{+}$. Figure 1 coincides with the well-known list of affine rigid isotopy classes of bounded degree four smooth algebraic plane curves. In our case, however, the isotopies are allowed only within the space $\Phi_{+}$ which is much smaller than the space of all degree four polynomials. In other cases, $\Phi_{1}$ and $\Phi_{2}$, the analogous restriction leads to an increase of classification lists.
@@
Using the Lyashko–Looijenga covering (in the form described in [12]), we can assume that the functions $f$ and $\tilde{f}$ have the same set of critical values and the same set of paths in ${\mathbb{C}}^{1}$ specifying the common associated virtual function. Let $I:[0,1]\to\Phi_{+}$ be a generic path such that $I(0)=f$, $I(1)$ is the polynomial ${\bf T}(x^{4}+x^{3}+y^{4})$, and the path $I$ has in its interior points only finitely many transversal intersections with the variety of non-generic polynomials in $\Phi_{+}$ (that is, with the union of the discriminant and the set of not strictly Morse functions). Since the virtual function of $\tilde{f}$ is the same as of $f$, there exists another path $\tilde{I}:[0,1]\to\Phi_{+}$ such that $\tilde{I}(0)=\tilde{f}$, the polynomials $I(t)$ and $\tilde{I}(t)$ have the same set of critical values for any $t\in[0,1]$, and the polynomials of the path $\tilde{I}$ undergo the surgeries of the same type as $I$ at the same instants $t$. (For this existence statement, the parabolicity of $X_{9}$ singularities is essential.) In particular, the point $\tilde{I}(1)$ can only be one of the four points (2). First, suppose that it is the same point $I(1)$. A neighborhood of this point in the space $\Phi_{+}$ has only one nontrivial smooth automorphism that preserves the associated virtual functions and commutes with the Lyashko–Looijenga map: this is the map induced by the reflection of ${\mathbb{R}}^{2}$ in the line $y=0$. Therefore, there are only two possibilities: either the final segments of the paths $I$ and $\tilde{I}$ (for $t\in(1-\varepsilon,1]$) are the same, or they are mapped into each other by this automorphism. In both cases, this condition can be continued to the entire paths $I$ and $\tilde{I}$ and proves that either $\tilde{f}=f$ or these two functions are mapped to each other by this automorphism, which belongs to the group $G_{0}$. If the point $\tilde{I}(1)$ is different from $I(1)$, we first apply the element of the group $G_{0}$, moving these points to each other, to the entire path $\tilde{I}$ (in particular, to its starting point $\tilde{f}$). This leads us to the situation considered previously. $\Box$
Proof of Theorem 6. By Proposition 9, any virtual component mentioned in Proposition 8 is associated with at most eight real connected components. Let us prove that in fact each of these seven virtual components is associated with a single real connected component (that is, these components are $G_{0}$-invariant).
@@
with arbitrary $A\in(-1,1)$. This polynomial is invariant under the transformations $(x,y)\mapsto(x,-y)$ and $(x,y)\mapsto(-x,y)$. Thus, its $G_{0}$-orbit consists only of it and the polynomial
@@
Proof. If we take $A=1$ in (5), then all these polynomials will be obtained from each other by the rotations of ${\mathbb{R}}^{2}$ by angles $\varphi\in[0,\pi/2]$. Each of these polynomials has only three real Morse critical points: one at the origin and two on the unit circle. After replacing $A=1$ by $1-\varepsilon$ with sufficiently small $\varepsilon$, the real zero-level set deforms only slightly in the domain $\{x^{2}+y^{2}\leq 2\}.$ In particular, it maintains its topological structure in this domain for any $\varphi$. It also remains empty outside this domain. $\Box$
@@ 2.4. The 1-cohomology of discriminant complements of $X_{9}^{+}$ singularities
Recall the notation $B({\mathbb{R}}^{2},k)$ for the unordered configuration space, that is, the space of $k$-point subsets of the plane. For any $k\geq 2$, $H^{1}(B({\mathbb{R}}^{2},k))\simeq{\mathbb{Z}}$. The basic element of this group is induced from $H^{1}({\mathbb{C}}^{1}\setminus\{0\})$ by the map $B({\mathbb{R}}^{2},k)\equiv B({\mathbb{C}}^{1},k)\to{\mathbb{C}}^{1}\{0\}$ sending any collection of $k$ distinct point $z_{1},\dots,z_{k}$ to the product of all k(k1)k(k-1) complex numbers $z_{i}-z_{j}$, $i\neq j$. Let $U$ be a connected component of the discriminant complement in $\Phi_{+}$ that consists of polynomials whose zero-level sets are unions of $k\geq 2$ non-nested ovals. Then a map $H^{*}(B({\mathbb{R}}^{2},k))\to H^{*}(U)$ is defined as follows. For any polynomial $f\in U,$ define $\beta(f)$ as the space of all choices of a point inside any of its ovals. These spaces $\beta(f)$ form a locally trivial fiber bundle over the component $U$. The fibers of this fiber bundle are contractible, hence it has a cross-section, which is a continuous map $U\to B({\mathbb{R}}^{2},k)$. Clearly, these maps defined by different cross-sections are homotopic and define the same homomorphism of cohomology rings.
@@ 2.5. The components for the $X_{9}^{1}$ singularities
acts on the space $\Phi_{1}$. It preserves the discriminant variety and hence permutes the connected components of its complement. This involution naturally extends to the set of associated virtual components: each virtual component which is the value of the set-valued invariant of the polynomial $f(x,y)$ goes to the value of the polynomial f(x,y)-f(x,-y). These involutions have no fixed elements: indeed, the sum of the values of the invariant Ind (see Definition 3) for any two components (real or virtual) related by the involution is always equal to 1-1.
@@
There are exactly ten virtual components of type $X_{9}^{1}$. They split into five pairs of virtual components related to each other by the involution ((13)). The topological types of zero-level sets of polynomials associated with elements of some five of the virtual components ((one from each pair)) are shown in pictures ((a)), ((b)), ((c)), ((d)), and ((f)) of Fig. 2. $\Box$
@@
Proof. Let $f$ be a polynomial that satisfies the conditions of this lemma. By Lemma 5 of [27], the $j$-invariant of its principal homogeneous part is equal to 0, hence $A=\pm\frac{\sqrt{3}}{2}$ in (12). We can assume that the coordinates $(x,y)$ are centered at the $E_{7}$ point. In these coordinates, the monomials of degree $\leq 2$ of the polynomial $f$ vanish, and the homogeneous part of degree three is proportional to the third degree of a linear function that also divides the degree four homogeneous part of $f$. Thus, the degree three homogeneous part has the form $\alpha x^{3}$ or $\alpha y^{3}$, $\alpha\neq 0$. The critical value of the polynomial at the $A_{2}$ point is equal to $\pm 3^{\frac{15}{2}}2^{-8}\alpha^{2}$, where $\pm$ is the sign of $A$. This value is equal to $3^{\frac{15}{2}}2^{-8}$ if and only if this sign is ++ and $\alpha$ equals $1$ or 1-1. $\Box$
Proof of Proposition 16. The deduction of this proof from Lemma 17 essentially repeats the proof of Proposition 9 with the following simplification. In this case, the neighborhood of the point ${\bf\tilde{T}}(x^{3}y+\sqrt{3}x^{2}y^{2}+xy^{3}+x^{3})$ in the parameter space of the deformation (1) has no nontrivial automorphisms that commute with the Lyashko–Looijenga covering (see [27], § 4.5). Therefore, the statement of Proposition 16 follows immediately from the fact that all four polynomials (14) belong to the same orbit of the group $G_{1}$. $\Box$
@@
Picture $(d)$ (respectively, $(e)$) is realized by the polynomials
(15) ${\bf\tilde{T}}\left((x^{2}+y^{2}-t^{2})(x+0.9t)(y+0.9t)+\alpha t^{4}\right),\quad\alpha\in(0,0.003),$
with positive (respectively, negative) $t$. These two pictures are topologically equivalent. However, it easily follows from Bezout’s theorem that these polynomials represent different connected components of the discriminant complement in the space $\Phi_{1},$ see [27].
@@ 2.6. Components for the $X_{9}^{2}$ singularities
There are exactly eight polynomials ${\mathbb{R}}^{2}\to{\mathbb{R}}$ of class $\Phi_{2}$ that have a critical point of class $D_{6}$ with critical value 0 and a critical point of class $A_{3}$ with critical value $\frac{-1}{324}$. These eight polynomials form an orbit of the natural action of the group $G_{0}$ on $\Phi_{2}$.
@@
The rest of the proof of Proposition 21 repeats the proof of Proposition 3 in § 5.4 of [27]. $\Box$
@@
whose zero-level set is topologically equivalent to the picture (d), but has a singular point of class $E_{6}$ at its center. Then, we perturb this singular point by adding very small terms of lower degrees. These terms do not change the topology of the curve far away from this singular point, but provide the standard dissolutions of the $E_{6}$ singularity, see Figure 24 in [30]. All the resulting polynomials (26)–(30) are invariant under the reflection in the axis $y=0$. Thus, in each of these five cases, the virtual component is associated with no more than four real connected components of the discriminant complement. These four components differ for the topological reasons.
@@ 3. Components for $J_{10}$ singularities
Denote by $\Omega_{1}$ and $\Omega_{3}$ the spaces of all polynomials of the form (31) and (32), respectively. The formal graph and its virtual components for $J_{10}$ singularities are defined in the same way as for the case $X_{9}$, except that the frames of vanishing cycles consist of ten elements.
@@
If two generic polynomials $f$, $\tilde{f}$ of class $\Omega_{1}$ or $\Omega_{3}$ are associated with the same virtual function, then their connected components of the space of generic polynomials either are the same or are mapped to each other by the involution
@@
In particular, any virtual component of class $\Omega_{1}$ or $\Omega_{3}$ is associated with at most two real connected components of the discriminant complements in $\Omega_{*}$. $\Box$
In the following pictures, it is assumed that the $x$ coordinate axis is directed upwards, and the horizontal line is the axis $\{x=0\}$.
@@ 3.2. Components for $J_{10}^{1}$ singularities
There are exactly ten virtual components of the formal graph of type $J_{10}^{1}$. These virtual components are realized by polynomials, the topological shapes of whose zero-level sets are shown on Fig. 4. $\Box$
@@
There are exactly thirteen connected components of the discriminant complement in the space of all polynomials of class $\Omega_{1}$. Each of the virtual components realized by polynomials with topological types shown in pictures $(a)$, $(b)$, (b~)($\tilde${b}), $(c)$, (c~)($\tilde${c}), $(d)$, and (d~)($\tilde${d}) of Fig. 4 is associated with a single real connected component. Each of the virtual components/topological types represented in pictures $(e)$, (e~),($\tilde${e}), and $(f)$ is associated with two different components that are related by the involution ((33)).
@@ 3.3. Components for $J_{10}^{3}$ singularities
(35) (g),(h),(h~),(i),(i~),(j),(j~),(k),(k~),(l)(g),\ (h),\ ($\tilde${h}),\ (i),\ ($\tilde${i}),\ (j),\ ($\tilde${j}),\ (k),\ ($\tilde${k}),\ (l)
@@ 4. The case of $P_{8}$ singularities
with $A<1$ for $P_{8}^{1}$ singularities and $A>1$ for $P_{8}^{2}$, see [5]. The zero sets of these normal forms of class $P_{8}^{1}$ (respectively, $P_{8}^{2}$) define cubic curves with one (respectively, two) connected component in ${\mathbb{R}P}^{2}$. Thus, each polynomial of class $P_{8}^{1}$ defines a conical surface in ${\mathbb{R}}^{3}$ that is homeomorphic to a plane and has a single singular point at the origin. The zero set of a polynomial of class $P_{8}^{2}$ consists of a surface like this and a surface that is homeomorphic to the standard quadratic cone.
@@
A polynomial $f:({\mathbb{C}}^{3},{\mathbb{R}}^{3})\to({\mathbb{C}},{\mathbb{R}})$ of the form (42) is generic if it has only Morse critical points in ${\mathbb{C}}^{3}$, its critical values are all distinct and not equal to 0, and the complex zero-level set of the principal homogeneous part of $f$ defines a smooth curve in ${\mathbb{C}}P^{2}$.
Let $\Xi_{1}$ and $\Xi_{2}$ denote the parameter spaces ${\mathbb{R}}^{7}\times(-\infty,1)$ and ${\mathbb{R}}^{7}\times(1,\infty)$ of the deformations (42) of the functions (41) of classes $P_{8}^{1}$ and $P_{8}^{2}$, respectively.